a korean restaurant menu contains 6 meat poultry, 4 stir fried dishes, 3 kimchi, 5 soups. how many meals are possible consisting 2 meats and poultry, 1 stir fy, 1 kimchi and 1 soup?
if the questions was 1 from each i would just 6*4*3*5=360
but since it's asking 2 meat poultry imma lil confused with whether if i should do
6*6*4*3*5 or 12*4*3*5?
thanks in a million.

Answer by stanbon(75887)

a korean restaurant menu contains 6 meat poultry, 4 stir fried dishes, 3 kimchi, 5 soups. how many meals are possible consisting 2 meats and poultry, 1 stir fy, 1 kimchi and 1 soup?
if the questions was 1 from each i would just 6*4*3*5=360
but since it's asking 2 meat poultry imma lil confused with whether if i should do 6*6*4*3*5 or 12*4*3*5?
thanks in a million.
-----------------------
2 meat/poultry in 6C2 = 15 ways
1 stir fry = 4 ways
1 kimchi = 3 ways
1 soup = 5 ways
---
Total = 15*4*3*5 = 900
===========================
Cheers,
Stan H.